OpenStudy (anonymous):

A wheel starts from rest and has an angular acceleration that is given by α(t) = (6rad/s^4)t^2 , The angle through which it turns in time t is given by ? I have got the answer 3t^4 , but the answer was (.5)t^4 !

OpenStudy (anonymous):

how did you get that? the answer is (.5)t^4... can you show me the work?

OpenStudy (anonymous):

i mean show me your work.. so ll tell where yu went wrong

OpenStudy (anonymous):

ok i got what you did.. and thats wrong.. you used the equation of motion.. that only works of the acceleration is a constant.. here.. the acceelration is clearly not a constant.. depends on time t.. hence you need to integrate!

OpenStudy (fifciol):

integrate twice your a(t) function. you will get: \[\int\limits_{}^{}6t^2dt=2t^3\] \[\int\limits_{}^{}2t^3dt=0.5t^4\]

OpenStudy (anonymous):

don't give out direct answers mate.. let him think ;-)!

OpenStudy (anonymous):

@Mashy ,, yes that right i made that mistake !

OpenStudy (anonymous):

yea so now you know right? :)

OpenStudy (fifciol):

ok @Mashy :)

OpenStudy (anonymous):

@Mashy ,, ,, yes I had known it :) thanx @Fifciol ,, Thank you too